# Physics-Informed Neural Networks for the Relativistic Burgers Equation in the Exterior of a Schwarzschild Black Hole

### v1

@@ 1 Introduction
Recent developments in machine learning and data analysis has led to results across multiple scientific domains. Among these, one of the most outstanding methods is physics informed neural networks (PINN) which incorporate physics laws and partial different equations directly in the training of the neural network. We are interested in exploring the application of PINNs to model compressible fluid flows in the exterior of a Schwarzschild black hole.
A key challenge is to take the curved spacetime geometry and the possible formation of shock waves into account when designing the network and the training process. The relativistic Burgers equation on a Schwarzschild background reads as
@@
where the unknown velocity $v\in[-1,1]$ with the normalized light speed and $M$ the Schwarzschild black hole mass. The flow is supposed to be spherically symmetric. It is straightforward to recover the standard Burgers equation when the mass $M$ vanishes.
@@
We organize the article as follows: In Section 2, we give a summary of the main results [2] as background on the relativistic Burgers equation Schwarzschild spacetime, including classical steady state and shock solutions. We also give an overview of Physics-Informed Neural Networks (PINNs) [3]. We introduce our relativistic PINNs Section 3. The model is designed to well capture shock formation by incorporating a shock-aware jump block and a Godunov-inspired residual construction. We also introduce the associated loss function and training strategy in detail. Section 4 presents numerical experiments and we demonstrate the efficiency s of our method in learning different type of solutions on a curved spacetime background. Finally, in Section 5, we discuss the pros and cos of our methods and the potential of PINNs as a mesh-free alternative for simulating relativistic fluid flows on complicated metrics.
@@ 2.1 Burgers Equation on Schwarzschild spacetime
or, equivalently, $s=(1-2M/r)\frac{v_{R}+v_{L}}{2}$ where $s$ is then speed of the shock. It is obvious that an initial condition with the form of (2.2) ensures a unique steady state solution of the same form. In addition to the smooth steady state solutions, we can define the class of steady shock solutions to the relativistic Burgers equation
@@ 2.2 Physics Informed Neural network
where $\mathcal{L}_{\text{PDE}}$ is the residual of the PDE at given collocation points, and $\mathcal{L}_{\text{BC}},\mathcal{L}_{\text{IC}}$ are boundary and initial conditions. PINNs have been successfully applied to many problems such as fluid mechanics, electromagnetism, etc. the
@@ 3.1 Model Formulation
    where $\sigma$ is the sigmoid function and $k$ is a trainable sharpness parameter. This acts as a differentiable indicator for the location of the shock.
@@ 3.2 Loss Function and Residual Construction
    $\mathcal{R}(t,r):=\frac{\partial_{t}v}{\text{metric}(r)^{2}}+\partial_{r}f(v).$ (3.3)

### v2

@@ 1 Introduction
Recent developments in machine learning and data analysis has led to results across multiple scientific domains. Among these, one of the most outstanding methods is physics informed neural networks (PINN) which incorporate physics laws and partial differential equations directly in the training of the neural network. We are interested in exploring the application of PINNs to model compressible fluid flows in the exterior of a Schwarzschild black hole.
In particular, we consider the relativistic Burgers equation on a Schwarzschild background reading as
@@
where the unknown velocity $v\in[-1,1]$ with the normalized light speed and $M$ the Schwarzschild black hole mass. We assume spherical symmetry for the flow. It is straightforward to recover the standard Burgers equation when the mass $M$ vanishes.
Previous studies on relativistic fluid dynamics in curved spacetimes, particularly by LeFloch et al. [1, 2], and on numerical methods by Glimm et al. [3, 4] and Russo et al. [11, 12], have primarily focused on developing discretization-based schemes to approximate shock wave solutions for the relativistic Burgers equation and the compressible Euler system. In contrast, our goal is to develop a mesh-free numerical approach using PINNs.
A key challenge is to take the curved spacetime geometry and the possible formation of shock waves into account when designing the network and the training process.
@@
We organize the article as follows:
In Section 2, we give a summary of the main results [9] as background on the relativistic Burgers equation Schwarzschild spacetime, including classical steady state and shock solutions. We also give an overview of Physics-Informed Neural Networks (PINNs) [10]. We introduce our relativistic PINNs Section 3. The model is designed to well capture shock formation by incorporating a shock-aware jump block and a Godunov-inspired residual construction. We also introduce the associated loss function and training strategy in detail. Section 4 presents numerical experiments and we demonstrate the efficiency s of our method in learning different type of solutions on a curved spacetime background. Finally, in Section 5, we discuss the pros and cos of our methods and the potential of PINNs as a mesh-free alternative for simulating relativistic fluid flows on complicated metrics.
@@ 2.1 Burgers Equation on Schwarzschild spacetime
or, equivalently, $s=(1-2M/r)\frac{v_{R}+v_{L}}{2}$ where $s$ is then speed of the shock. It is obvious that an initial condition with the form of (2.2) ensures a unique steady state solution of the same form. In addition to the smooth steady state solutions, we can define a class of steady shock solutions to the relativistic Burgers equation
@@ 2.2 Physics Informed Neural network
where $\mathcal{L}_{\text{PDE}}$ is the residual of the PDE at given collocation points, and $\mathcal{L}_{\text{BC}},\mathcal{L}_{\text{IC}}$ are boundary and initial conditions. PINNs have been successfully applied to many problems such as fluid mechanics, electromagnetism, etc.
Extensions and variations have been proposed ever since: Adaptive PINNs (APINNs) improve convergence by dynamically re-weighting different loss components, handling especially regions with steep gradients [5]; Extended PINNs(XPINN) applies a domain decomposition and adopt multiple neural networks to solve PDEs in subdomains and helps the case where the geometry is complicated [6]; Conservative PINNs(cPINNs) introduce the conservation laws into the model and this is of essential importance for hyperbolic systems [7].
@@ 3.1 Model Formulation
    where $\sigma$ is the sigmoid function and $k$ is a trainable sharpness parameter which controls the steepness of the sigmoid transition.This acts as a differentiable indicator for the location of the shock.
@@ 3.2 Loss Function and Residual Construction
    $\mathcal{L}_{\text{loss}}(t,r):=\frac{\partial_{t}v}{\text{metric}(r)^{2}}+\partial_{r}f(v).$ (3.3)